OpenStudy (anonymous):

Find the slope of the line containing the given points. (2, 3), (5, 6)

OpenStudy (anonymous):

Do you know the formula for the slope of a line?

OpenStudy (amistre64):

the steps work better for me

OpenStudy (amistre64):

subtract one point from the other and stack the ys/xs

OpenStudy (anonymous):

@jabberwock i believe its rise over run . .

OpenStudy (anonymous):

Right. Which coordinate represents the rise?

OpenStudy (anonymous):

*x or y*

OpenStudy (anonymous):

x

OpenStudy (anonymous):

x goes left and right, and y goes up and down. So the "rise" refers to a change in the y-coordinates, and the "run" refers to a change in the x-coordinates.

OpenStudy (anonymous):

What are your two y-coordinates?

OpenStudy (anonymous):

3,6

OpenStudy (anonymous):

Good, so subtract one of those from the other.

OpenStudy (anonymous):

3-6=?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

Good :)

OpenStudy (anonymous):

-3 right

OpenStudy (anonymous):

Now the x-coordinates. The only thing you have to keep in mind in finding the slope is that you need to keep the coordinates in the same order. The first y-coordinate you used was 3, which was part of the point (2, 3). This means you need to calculate 2-5 (instead of 5-2)

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

So you have -3/-3, which is equal to +1

OpenStudy (anonymous):

rise/run

OpenStudy (anonymous):

Or would it be easier if i use the formula \[y-y_1=m(x-x_1)\]

OpenStudy (precal):